OpenStudy (anonymous):

Calib received $50 for his birthday and put it in his piggy bank. Each week, he puts 2 more dollars in his bank. The amount of money in dollars D in his bank is given by D = 2w + 50 where w is the number of weeks he has saved 2 dollars. His sister Lara makes a table of reasonable values for w. Lara selects 14 as a value for w. What is the corresponding table entry for D?

OpenStudy (anonymous):

d=2(14)+50 = 78, The reason I kept 14 was because the number of weeks the dollars were saved x the value that is 14

OpenStudy (anonymous):

D = 78 D= 2(14) You substitute 14 for "w" because thats the number that was given to you. D= 28 , Because 2x14 = 28 D= 28+50 which = 78